In Country B, assume that banks have no excess reserves to start with and cash drain is 50%. Assume a required reserve ratio of 10%. The Central Bank of Country B decreases the monetary base (i.e. high-powered money) by 100,000 Liras. a) Explain how this policy will affect the money supply. Calculate the change in money supply and explain how the change comes about. b) How can the Central Bank decrease the monetary base by 100,000 Liras? c) Suggest an alternative tool that the Central Bank can employ to bring about the same change in the money supply as calculated in (a). d) What policy objectives could the Central Bank be pursuing so as to want to decrease the monetary base? e) What are some limitations of the Central Bank in achieving the objectives you mention in (d)?
1
Expert's answer
2020-06-10T19:30:19-0400

a)

Money supply decreases and money supply decreases

"m=\frac{1}{reserve ratio}" "m=\frac{1}{reserve ratio}"


"M=m\times B"

"m=\frac{1}{0.1}=10"

Let the original initial amount was 200,000


"M=10\times200 000=2 000 000"

after decreasing

"M=10\times100 000=1 000 000"

b)

Central Bank may increase reserve requirement

c)

Refinancing Rate Change, Discounting, rediscounting,

foreign exchange intervention

d)

Decrease in inflation in the country, cooling of an overheated economy

e)

The Central Bank cannot constantly lower the refinancing rate, raise the reserve ratio, etc. This is a regulatory body, therefore, analyzing the state of the economy indicated in paragraph d, the Central Bank changes the above parameters